the perimeter of a rectangle is 84 cm. The lenght is twice the width. find the length and width.




Answer by bam878s(77)

The permiter is simply the distance around the object. For example, if you wish to fence in a yard, you would calculate the distance around the yard (the permieter) in order to find the amount of fencing material you will need.
If we let a = length and b = width, then we know that the perimeter is 2a+2b = 84. Since the length is twice the width we also know that a = 2b. So now we simply plug this into our orignal formula.
2a + 2b = 2(2b) + 2b = 4b + 2b = 84.
Now, we can solve for b.
4b + 2b = 84
6b = 84
b = 14
Now we know length ("a") is twice the width ("b"). So,
a = 2b = 2(14) = 28
So a=28 and b=14 so this is your length and width respectively.
Hope this helps!